# Vision-based Learning for Drones: A Survey

### 08 Dec 2023 v1

@@ 7 Conclusion
This comprehensive survey has thoroughly explored the rapidly growing field of vision-based learning for drones, particularly emphasizing their evolving role in multi-drone systems and complex environments such as search and rescue missions and adversarial settings. The investigation revealed that drones are increasingly becoming sophisticated, autonomous systems capable of intricate tasks, largely driven by advancements in AI, machine learning, and sensor technology. The exploration of micro and nano drones, innovative structural designs, and enhanced autonomy stands out as key trends shaping the future of drone technology. Crucially, the integration of visual perception with machine learning algorithms, including deep reinforcement learning, opens up new avenues for drones to operate with greater efficiency and intelligence. These capabilities are particularly pertinent in the context of object detection and decision-making processes, vital for complex drone operations. The survey categorized vision-based control methods into indirect, semi-direct, and end-to-end methods, offering a nuanced understanding of how drones perceive and interact with their environment. Applications of vision-based learning in drones, spanning from single-agent to multi-agent and heterogeneous systems, demonstrate their versatility and potential in various sectors, including agriculture, industrial inspection, and emergency response. However, this expansion also brings forth challenges such as data processing limitations, real-time decision-making, and ensuring robustness in diverse operational scenarios.

### 02 Jan 2024 v2

@@ 7 Conclusion
This comprehensive survey has thoroughly explored the rapidly growing field of vision-based learning for drones, particularly emphasizing their evolving role in multi-drone systems and complex environments such as search and rescue missions and adversarial settings. The investigation revealed that drones are increasingly becoming sophisticated, autonomous systems capable of intricate tasks, largely driven by advancements in AI, machine learning, and sensor technology. The exploration of micro and nano drones, innovative structural designs, and enhanced autonomy stand out as key trends shaping the future of drone technology. Crucially, the integration of visual perception with machine learning algorithms, including deep reinforcement learning, opens up new avenues for drones to operate with greater efficiency and intelligence. These capabilities are particularly pertinent in the context of object detection and decision-making processes, vital for complex drone operations. The survey categorized vision-based control methods into indirect, semi-direct, and end-to-end methods, offering a nuanced understanding of how drones perceive and interact with their environment. Applications of vision-based learning drones, spanning from single-agent to multi-agent and heterogeneous systems, demonstrate their versatility and potential in various sectors, including agriculture, industrial inspection, and emergency response. However, this expansion also brings forth challenges such as data processing limitations, real-time decision-making, and ensuring robustness in diverse operational scenarios.